What is the formal charges on atoms: CSO? What is the hybridization ion on central atom?
Answer #1

CSO :S= C= 0 FC on C = 4 - 4 -0 = 0 FC on S = 6-2 - 4 = 0 FC on 0 = 6-2-4 = 0 hybridization of C = sp